Can you help me solve this I need to show work also ? 9g+2+2g+3=82 (Math is my hardest subject)
Answer by mananth(16946)

9g+2+2g+3=82
add the like terms
9g+2g+2+3=82
11g+5=82
11g+5-5=82-5
11g=77
divide by 11
11g/11 = 77/11
g=7